Please answer all parts of the question: a,b,c,dIdeal Gas Law The ideal gas law states that PV = Nk T where P is the absolute pressure of a gas, V is the volume it occupies,
Answer #1

(a) Just use ideal gas law

PV = nRT

n = PV / RT

where R is gas constant

n = 7e5 * 2e-3 / 8.314 * (273 + 26)

n = 0.563

-------------------

(b) Pf / Po = Tf / To

Pf = (Tf / To ) * Po

Pf = ( 35 + 273 / 26 + 273) * 7e5

Pf = 7.21e5 Pa

(write only 7.21 in the box)

--------------------------

(c)

here volume is kept constant

so,

T = 3 * T1

T = 3 * (273 + 26)

T = 897 K

or

T = 624 oC

-------------------------------------

(d)

Vf = ( 1.5 / 299) * 348

Vf = 1.745 m3